3) A 650-N billboard worker stands on a 4.00-m scaffold weighing 600 N and supported by vertical ropes at each end. How far would the worker stand from one of the supporting ropes to produce a tension of 660 N in that rope and 590 in the other rope?

a. 1.78 m

d. 0.06 m

b. 2.00 m

c. 1.90 m

e. 0.93 m
Answer #1

660N 590 other and that end 650N GOON

Using Net torque balance about that rope from above FBD:

Suppose worker is at 'x' distance from that rope, then

Clockwise torque = Counterclockwise torque

590*4 = 650*x + 600*(4.0/2)

x = (590*4 - 600*2)/650

x = 1.78 m from that end

Correct option is A.

Please Upvote.